How can I find the solution to this problem? "Assume that is an angle in standard position whose terminal side contains the point (-24, 7). Find the exact values for the trig functions"
Answer by lwsshak3(6490)

How can I find the solution to this problem? "Assume that is an angle in standard position whose terminal side contains the point (-24, 7). Find the exact values for the trig functions".
..
Given info places the reference angle (x) in quadrant II.
You will be working with a right triangle with a vertical leg=7 and a horizontal leg=-24.
This makes the hypotenuse=sqrt(24^2+7^2)=sqrt(625)=25.
..
sin x=7/25
cos x=-24/25
tan x=-7/24
csc x=25/7
sec x=-25/24
cot x=-24/7